OpenStudy (anonymous):

find range of sqrt(25-x^2-y^2-z^2)

OpenStudy (anonymous):

\[\sqrt{25-x^{2}-y^{2}-z^{2}}\]

OpenStudy (anonymous):

25 - x^2 - y^2 - z^2 ≥ 0 25 ≥ x^2 + y^2 + z^2 5^2 ≥ x^2 + y^2 + z^2 so it's a sphere with the radius less than or equal to 5

OpenStudy (anonymous):

actually it's the upper hemisphere

OpenStudy (anonymous):

ok, just use sphere formula to find the range?? can't be calculate by using function way??

OpenStudy (anonymous):

actually this is the multivariable function question

OpenStudy (anonymous):

f(x,y,z)=sqrt(25-x^2-y^2-z^2)

OpenStudy (anonymous):

if that's the case, 0 <= f(x,y,z) <= 5

OpenStudy (anonymous):

ya, but how to calculate??

OpenStudy (anonymous):

still using sphere formula??

OpenStudy (anonymous):

well it's not sphere anymore. It's like a 4-dimension "sphere"

OpenStudy (anonymous):

quite blur.... is there any method of calculation??

OpenStudy (anonymous):

well, i'm certian there are.

OpenStudy (anonymous):

sqrt(r^2 - x^2) is a upper semi-circle and the range is 0 <= y = < r sqrt(r^2 - x^2 - y^2) is an upper hemisphere, range: 0 <= z <= r sqrt(r^2 - x^2 - y^2 - z^2) (this is 4-d), regardless, range: 0 <= f(x,y,z) <= r

OpenStudy (anonymous):

ok. thx a lot, understand ald... ^^